edmund purchased 8/9 acre of land in one year and 5/9 acre the next year. he planted rice 7/9 acre of the land and corn on the remainder. How much land is planted for corn?


1
Expert's answer
2021-11-09T01:07:44-0500

total land purchase = "\frac{8}{9}+\frac{5}{9}=\frac{13}{9}" arc

rice are planted = "\frac{7}{9}" arc

corn planted = total land purchase - rice are planted

="\frac{13}{9}-\frac{7}{9}"

="\frac{6}{9}\\\frac{2}{3}"